Night Logs from 2010

28-30 Dec 2010 UT LUCIFER/LBC Science Prep (combined, mostly bad weather).
01 Nov 2010 UT LUCIFER/LBC Science Prep.
31 Oct 2010 UT LUCIFER/LBC Science Prep.
29 Sep 2010 UT LUCIFER Final check before science run.
20 Sep 2010 UT LBC Restart after shutdown - We seem to have removed ~50 micron tilt on LBC Blue. Closed just before 10pm for high humidity. Still would like to re-check rotator center and double-check tilt of LBCB, check filter focus offsets.
19 Sep 2010 UT LBC Restart after shutdown - Obtained Pointing Model data before midnight, After midnight (poor seeing) tested Master/Slave guiding and automatic tech chip focusing.
18 Sep 2010 UT LBC Restart after shutdown - Took data to assess tilt, measure rotator center, construct new collimation models)
17 Sep 2010 UT LBC Restart after shutdown - Diagnosed problem with chip 2 is with video cable
14 Sep 2010 UT LFBG/LUCIFER Restart after shutdown

Duplicate LUCIFER Mask Checks

02 Jul 2010 UT LUCIFER Training Obs
01 Jul 2010 UT LUCIFER Training Obs

06 June 2010 UT LUCIFER & LBC Science Prep
05 June 2010 UT LUCIFER & LBC Science Prep
04 May 2010 UT LUCIFER Science Prep & Training
03 May 2010 UT - lost to bad weather
02 May 2010 UT LUCIFER Science Prep & Training
01 May 2010 UT LUCIFER Science Prep & Training
08 April 2010 UT LBC,LUCIFER Science Prep
07 April 2010 UT LBC,LUCIFER Science Prep

-- DavidThompson - 21 Feb 2011
Topic revision: r1 - 21 Feb 2011, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback